OpenStudy (anonymous):

Put the equation in the form ax^2+bc+c=0 (2x+1)^2-3x=5

OpenStudy (anonymous):

let me try kay

OpenStudy (anonymous):

2x^2-3x+5=0

OpenStudy (anonymous):

YOU BETTER TRY IT :P \[(2x+1)^2-3x=5 \]\[4 x^2+x+1 = 5\] You finish...

OpenStudy (anonymous):

damnn i suck at this xD

OpenStudy (anonymous):

did you figure it out?